OpenStudy (anonymous):

Luke has a spinner divided into 5 equal segments, each with a different color. In fifty spins, he spun yellow twenty times. How does this compare to the expected result? The experiment resulted in ten more yellows than expected. The experiment resulted in twenty more yellows than expected. The experiment resulted in ten fewer yellows than expected. The experiment resulted in twenty fewer yellows than expected.

OpenStudy (anonymous):

The experiment resulted in ten more yellows than expected

OpenStudy (anonymous):

Possible outcome: 5 colours Probability of yellow = 1 / 5 Number of yellow expected = Probability of yellow x number of spin = 1/5 x 50 = 10. Obtained result : 20 Expected result : 10

OpenStudy (anonymous):

This is because we spin 50 times with 5 different but equal color segments. 50/5=10 expected yellow spins. We have 20, and therefore 10 more than expected yellow spins.